stomach relief medications contains the compound Sodium hydrogen carbonate; an over the counter bottle of the medication contains 350g of Sodium hydrogen carbonate.
Calculate the number of particles in this sample of Sodium hydrogen carbonate.
Calculate the number of oxygen ions present in this mass.
1
Expert's answer
2020-10-14T06:00:07-0400

The number of the particles of the sodium hydrogen carbonate can be calculated from the number of the moles "n" of NaHCO3 using the Avogardro's number "N_A = 6.022\cdot10^{23}" mol-1:

"N = n\cdot N_A" .

The number of the moles of NaHCO3 equals its mass "m=350" g, divided by its molar mass "M = 84.01" g/mol:

"n = \frac{m}{M} = \frac{350 \text{ g}}{84.01\text{ g/mol}} = 4.17" mol.

Thus:

"N = 4.17\text{ mol} \cdot 6.022\cdot10^{23}\text{ mol}^{-1} = 2.51\cdot10^{24}" .

In every NaHCO3 unit, there are 3 ions of oxygen. Therefore, the number of oxygen ions present in this mass is three times the number of NaHCO3 molecules:

"N(O) = 3N = 7.53\cdot10^{24}" .

Answer: In 350 g of sodium hydrogen carbonate, there are 2.51x1024 molecules of NaHCO3 and 7.53x1024 ions of oxygen.